OpenStudy (anonymous):

Elliot and Katy both bought the same lunchbox for $6. Elliot lives in Oklahoma and pays 4.5% in sales tax, while Katy lives in South Carolina and pays 6% in sales tax. How much more did Katy pay in sales tax than Elliot? A.$0.09 B.$0.27 C.$0.63 D.$0.36

OpenStudy (ganpat):

use the same formula as studied before, separately for each of them and than find the difference :D

OpenStudy (anonymous):

i dont like this one tho -_____-

OpenStudy (anonymous):

i say C

OpenStudy (ganpat):

i doubt that...

OpenStudy (ganpat):

can u give one more try over here.. so you can understand, where u r going wrong..

OpenStudy (ganpat):

I know u dont like, but it has its own importance.. thats y homework is give.. so that u practice.. and be perfect...

OpenStudy (ganpat):

elliot = 6 * 4.5 /100 + 6 = 6.27 Katy = 6 * 6 /100 + 6 = 6.36 so, 6.36 - 6. 27 = 0.09 $... its C